OpenStudy (anonymous):

Find the value of squared 8 over 3 squared 4. Please help quick, need answer asap!

OpenStudy (anonymous):

64/(3sqrt(4)) 10.667

OpenStudy (anonymous):

OH OOPS JUST KIDDING THAT WAS SQUARED 4.

OpenStudy (whpalmer4):

jesusisabiscuit, that doesn't look right — you've taken a square root in one place and a square in another, but in both places the original problem says "squared"

OpenStudy (anonymous):

1.333

OpenStudy (anonymous):

A- 6 B- 2/3 C- sqrt 2/3 D- 3/sqrt 2

OpenStudy (whpalmer4):

zink96, the problem is actually \[\sqrt{8}/(3\sqrt{4})\], right?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Yeah, that's "square root", not "squared"! 4= 2*2 and 8 = 2*2*2, so the top simplifies to \[2\sqrt{2}\] and the bottom simplifies to \[3*2\] leaving \[\sqrt{2}/3\] after canceling out the matching numbers on top and bottom.

OpenStudy (anonymous):

Oh sorry thats what I meant. Thanks